The length of a rectangle is 2 cm longer than its width. If the diagonal of the rectangle is 4 cm, what are the dimensions of the rectangle?
Answer by jorel1380(3719)

Let w be the width of the triangle. Then the length of the triangle is w+2. So:
w²+(w+2)²=4²=16
w²+w²+4w+4=16
2w²+4w-12=0
w²+2w-6=0
(w+1)²=7
w+1=√7
w=√7-1
w+2=√7+1
The dimensions are √7-1 by √7+1
☺☺☺☺